Circular No. 6182

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995R IN UGC 8801 R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of a supernova (mag ∼ 18) on a deep U.K. Schmidt telescope R survey film taken by Q. A. Parker on June 29.5 UT; a similar film taken by M. J. Drinkwater on June 2 does not show the object to mag 21. A confirming unfiltered CCD image was obtained by G. J. Garradd and McNaught with the 1.0-m reflector at Siding Spring on July 6. SN 1995R is located at α = 13^{h}53^{m}32^{s}.60, δ = −1^{o}11′30′′.2 (equinox 2000.0), which is 9′′ east and 10′′ north of the center of UGC 8801 = MCG +00-35-025. A nearby star of mag 17 has position end figures 32^{s}.24, 31′′.3. S. Benetti, European Southern Observatory (ESO), writes: “I have observed SN 1995R on July 6.03 UT with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph) at La Silla. The partially-reduced spectrum (range 356-749 nm, resolution 0.4 nm) resembles those of a type-Ia supernova, ∼ 3 weeks past maximum light. Broad emissions are measured at 472, 504, 541.5, 554, 602, and 659.5 nm. The Si II 635.5-nm absorption is measured at 630.2 nm.”

GRO J1629–49 G. Kanbach, Max-Planck-Institut für Extraterrestrische Physik, and

the EGRET instrument team communicate: “A new, very bright γ-ray source has been detected in an EGRET quick-look analysis of the current pointing of the Compton Gamma Ray Observatory. The preliminary galactic position of GRO J1629–49 is l = 335^{o}.0, b = −0^{o}.4 (95-percent confidence radius ∼ 0^{o}.7). At energies above 100 MeV, this is one of the brightest sources ever detected. Previous exposures of this region have never shown a significant source at this location. The Parkes-Molonglo-NRAO catalog of radio sources contains an unresolved source, PMN J1628–4936, at α = 16^{h}28^{m}57^{s}.7, δ = −49^{o}36′25′′ (equinox 2000.0), with flux density 1134 ± 60 mJy at 4850 MHz during June 1990. PKSCAT90 does not list this source (2.7 GHz), which could indicate a time-variable radio emission. The other radio sources in the γ-ray error box are either extended or much weaker than the above object. Immediate multi-wavelength observations of PMN J1628– 4936 and the γ-ray uncertainty box of GRO J1629–49 are of strong interest. The current pointing of the Compton observatory will be maintained until July 10. Please contact G. Kanbach (gok@mpe-garching.mpg.de) if you can provide data on this source.”

1995 July 6 Daniel W. E. Green